OpenStudy (anonymous):

Which sum or difference identity would you use to verify that sin (90° + (theta) = cos (Theta)? a. cos (a -b) = cos a cos b + sin a sin b b. sin (a +b) = sin a cos b + cos a sin b c. sin (a -b) = sin a cos b- cos a sin b d. cos (a +b) = cos a cos b - sin a sin b

OpenStudy (alekos):

Well. I reckon it would have to be either b or c. What do you think?

OpenStudy (anonymous):

It looks to be that they are asking for the sum of identities. I think, not entirely sure, but I think... So this brings me to believe it, B.

OpenStudy (anonymous):

its c, i think so

OpenStudy (anonymous):

@Krishtika Well, looking at the equation I see this. (90° + (theta)) which makes me think a+b and, I know that it is not d.

zepdrix (zepdrix):

\[\Large\rm \sin(90^o+\theta)=\sin90^o \cos \theta+ \cos 90^o \sin \theta\] \[\Large\rm =(1) \cos \theta+ (0) \sin \theta\] Mmm ya it looks using the `sum` formula gives us the correct result, ya? :)

OpenStudy (anonymous):

@zepdrix So, Am I correct with it being B?

zepdrix (zepdrix):

yes, good job \c:/

OpenStudy (anonymous):

Thank you @zepdrix